INFRARED SPECTROSCOPY OF SIZE-SELECTED PROTONATED MOLECULAR CLUSTERS

\underline{G. E. DOUBERLY}, B. W. TICKNOR, A. M. RICKS, and M. A. DUNCAN, Department of Chemistry, University of Georgia, Athens, GA 30602.

Protonated molecular clusters such as (H_{2}O)_{n}H^{+}, (CO_{2})_{n}H^{+}, and (C_{2}H_{2})_{n}H^{+} are formed in a pulsed electric discharge of a supersonic expansion. The ions are mass selected and fragmented upon resonant vibrational excitation with the output of a tunable infrared OPO. The infrared spectra of these species provide valuble insight into the dynamics of the solvated proton.